# Stochastic Optimization with Random Search

### v1

@@ C.2 Analysis under individual Smoothness
$\mathbb{E}\big[f({\bm{x}}_{t+1})\big]$ $\leq\mathbb{E}\big[f({\bm{x}}_{t})\big]-\tfrac{\mu_{\mathcal{D}}}{2}\,\eta\,\mathbb{E}\big\|\nabla f({\bm{x}}_{t})\big\|+\tfrac{\mathcal{L}_{0}+\mathcal{L}_{1}\sigma}{2}\,\eta^{2}+\frac{4\eta}{\sqrt{db}}\,\sigma+\frac{8\sqrt{2}\,\eta^{2}}{\sqrt{b}}\Big(\mathcal{L}_{0}+\mathcal{L}_{1}\sigma+\mathcal{L}_{1}\,\mathbb{E}\big\|\nabla f({\bm{x}}_{t})\big\|\Big)$
$=\mathbb{E}\big[f({\bm{x}}_{t})\big]-\Big(\tfrac{\mu_{\mathcal{D}}}{2}\,\eta-\tfrac{8\sqrt{2}\,\mathcal{L}_{1}}{\sqrt{b}}\,\eta^{2}\Big)\mathbb{E}\big\|\nabla f({\bm{x}}_{t})\big\|+\underbrace{\Big(\tfrac{\mathcal{L}_{0}+\mathcal{L}_{1}\sigma}{2}\,\eta^{2}+\tfrac{4\eta}{\sqrt{db}}\,\sigma+\tfrac{8\sqrt{2}\,\eta^{2}}{\sqrt{b}}(\mathcal{L}_{0}+\mathcal{L}_{1}\sigma)\Big)}_{=:~\beta(\eta,b)}.$
Let $\alpha(\eta,b)\;:=\;\tfrac{\mu_{\mathcal{D}}}{2}\,\eta-\tfrac{8\sqrt{2}\,\mathcal{L}_{1}}{\sqrt{b}}\,\eta^{2}.$ If we choose:
$\eta\;\leq\;\min\!\left\{\frac{\mu_{\mathcal{D}}}{\mathcal{L}_{1}},\,\frac{\mu_{\mathcal{D}}\sqrt{b}}{32\sqrt{2}\,\mathcal{L}_{1}}\right\},$
then $\alpha(\eta,b)\geq\tfrac{\mu_{\mathcal{D}}}{4}\,\eta$. Summing the one–step bound over $t=0,\ldots,T-1$ and telescoping gives:
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big\|\nabla f({\bm{x}}_{t})\big\|\;\leq\;\frac{4F_{0}}{\mu_{\mathcal{D}}\eta\,T}\;+\;\frac{2\mathcal{L}_{0}+2\mathcal{L}_{1}\sigma}{\mu_{\mathcal{D}}}\,\eta\;+\;\frac{16}{\mu_{\mathcal{D}}}\,\frac{\sigma}{\sqrt{db}}\;+\;\frac{32\sqrt{2}}{\mu_{\mathcal{D}}}\,\frac{\eta}{\sqrt{b}}\,(\mathcal{L}_{0}+\mathcal{L}_{1}\sigma).$
@@
We choose the stepsize $\eta$ minimizing the right-hand side of Theorem C.3 subject to $\eta\leq\min\!\left\{\tfrac{\mu_{{\mathcal{D}}}}{{\mathcal{L}}_{1}},\,\tfrac{\mu_{{\mathcal{D}}}\sqrt{b}}{32\sqrt{2}\,{\mathcal{L}}_{1}}\right\}$, namely
$\eta\;=\;\min\!\left(\tfrac{\mu_{{\mathcal{D}}}}{{\mathcal{L}}_{1}},\,\tfrac{\mu_{{\mathcal{D}}}\sqrt{b}}{32\sqrt{2}\,{\mathcal{L}}_{1}},\,\sqrt{\tfrac{F_{0}}{({\mathcal{L}}_{0}+{\mathcal{L}}_{1}\sigma)\,T}}\right).$ (25)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla f({\bm{x}}_{t})\|\;=\;\frac{1}{\mu_{{\mathcal{D}}}}\,{\mathcal{O}}\!\left(\frac{{\mathcal{L}}_{1}F_{0}}{\mu_{{\mathcal{D}}}T}+\sqrt{\tfrac{({\mathcal{L}}_{0}+{\mathcal{L}}_{1}\sigma)\,F_{0}}{T}}+\frac{\sigma}{\sqrt{db}}\right).$ (26)
@@
For the uniform distribution over the unit sphere, we have: $\mu_{{\mathcal{D}}}\asymp d^{-1/2}$. With $T=\Theta\!\big(d\mathcal{L}_{1}/\varepsilon+d(\mathcal{L}_{0}+\sigma\mathcal{L}_{1})F_{0}/\varepsilon^{2}\big)$ and $b=\Theta(\sigma^{2}/\varepsilon^{2})$ we can ensure that $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla f({\bm{x}}_{t})\|\leq\epsilon$ , so the total number of function evaluations to ensure that $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla f({\bm{x}}_{t})\|\leq\epsilon$, is
$\textsc{Calls}=2bT\;=\;{{\mathcal{O}}}\!\left(\frac{d(\mathcal{L}_{0}+\mathcal{L}_{1}\sigma)F_{0}\,\sigma^{2}}{\varepsilon^{4}}\right).$
@@ Appendix D Finite sum case and Variance reduction
Under Assumption equation 27, suppose that $f$ is $(L_{0},L_{1})$-smooth, ${\bm{s}}_{t}\sim{\mathcal{D}}$ satisfies Assumption 2, and a variance-reduction scheme with snapshot period $m$ is used. Let $\eta_{t}\equiv\eta\leq\mu_{\mathcal{D}}/L_{1}$ and $F_{0}:=f({\bm{x}}_{0})-f^{\star}<\infty$. Then for each iteration,
$\mathbb{E}[f({\bm{x}}_{t+1})]\;\leq\;\mathbb{E}[f({\bm{x}}_{t})]\;-\;\frac{\mu_{{\mathcal{D}}}}{2}\,\eta\,\mathbb{E}\|\nabla f({\bm{x}}_{t})\|\;+\;\frac{L_{0}}{2}\,\eta^{2}\;+\;\frac{4\,G\,m}{\sqrt{b}}.$ (28)
Averaging over $t=0,\dots,T-1$, we obtain
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla f({\bm{x}}_{t})\|\;\leq\;\frac{1}{\mu_{{\mathcal{D}}}}\left(\frac{F_{0}}{\eta T}+\frac{L_{0}}{2}\,\eta+\frac{4\,G\,m}{\sqrt{b}}\right).$ (29)
@@ Appendix E Learning with Helper or Human Feedback: Convergence Proof
Suppose $f$ is $(L_{0},L_{1})$-smooth in the sense of equation 13, the random directions satisfy Assumption 2 ($\mathbb{E}\|{\bm{s}}_{t}\|^{2}=1$ and $\mathbb{E}[|\langle\nabla f({\bm{x}}),{\bm{s}}_{t}\rangle|]\geq\mu_{{\mathcal{D}}}\|\nabla f({\bm{x}})\|$), and the helper satisfies Assumption 8. Let $F_{0}:=f({\bm{x}}_{0})-f^{\star}<\infty$ and choose a constant stepsize $\eta_{t}\equiv\eta\leq\mu_{{\mathcal{D}}}/L_{1}$. Then
$\mathbb{E}[f({\bm{x}}_{t+1})]\;\leq\;\mathbb{E}[f({\bm{x}}_{t})]\;-\;\frac{\mu_{{\mathcal{D}}}}{2}\,\eta\,\mathbb{E}\|\nabla f({\bm{x}}_{t})\|\;+\;\frac{L_{0}}{2}\,\eta^{2}\;+\;2\,\delta.$ (34)
Averaging over $t=0,\dots,T-1$ yields
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla f({\bm{x}}_{t})\|\;\leq\;\frac{1}{\mu_{{\mathcal{D}}}}\left(\frac{F_{0}}{\eta T}+\frac{L_{0}}{2}\,\eta+\frac{2\delta}{\eta}\right).$ (35)
@@ Appendix F Failure of Classical Momenta
We now examine whether classical momentum techniques from first-order optimization can improve the performance of Random Search. We consider three natural extensions of popular momentum formulations: Heavy-Ball [Polyak, 1964], momentum-based variance reduction (MVR) [Cutkosky and Orabona, 2019], and implicit gradient transport [Arnold et al., 2019]. Each method can be adapted to our comparison-based framework by replacing stochastic gradients with function differences.

### v2

@@ C.2 Analysis under individual Smoothness
$\mathbb{E}\big[f({\bm{x}}_{t+1})\big]$ $\leq\mathbb{E}\big[f({\bm{x}}_{t})\big]-\tfrac{\mu_{\mathcal{D}}}{2}\,\eta\,\mathbb{E}\big\|\nabla f({\bm{x}}_{t})\big\|+\tfrac{L_{0}+L_{1}\sigma}{2}\,\eta^{2}+\frac{4\eta}{\sqrt{db}}\,\sigma+4\eta^{2}\Big(L_{0}+L_{1}\sigma+L_{1}\,\mathbb{E}\big\|\nabla f({\bm{x}}_{t})\big\|\Big)$
$=\mathbb{E}\big[f({\bm{x}}_{t})\big]-\Big(\tfrac{\mu_{\mathcal{D}}}{2}\,\eta-4\eta^{2}L_{1}\Big)\mathbb{E}\big\|\nabla f({\bm{x}}_{t})\big\|+\underbrace{\Big(\tfrac{L_{0}+L_{1}\sigma}{2}\,\eta^{2}+\tfrac{4\eta}{\sqrt{db}}\,\sigma+4\eta^{2}(L_{0}+L_{1}\sigma)\Big)}_{=:~\beta(\eta,b)}.$
Let $\alpha(\eta,b)\;:=\;\tfrac{\mu_{\mathcal{D}}}{2}\,\eta-4\eta^{2}L_{1}.$ If we choose: $\eta\;\leq\;\frac{\mu_{\mathcal{D}}}{16L_{1}}$, then $\alpha(\eta,b)\geq\tfrac{\mu_{\mathcal{D}}}{4}\,\eta$. Summing the one–step bound over $t=0,\ldots,T-1$ and telescoping gives:
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big\|\nabla f({\bm{x}}_{t})\big\|\;\leq\;\frac{4F_{0}}{\mu_{\mathcal{D}}\eta\,T}\;+\;\frac{2L_{0}+2L_{1}\sigma}{\mu_{\mathcal{D}}}\,\eta\;+\;\frac{16}{\mu_{\mathcal{D}}}\,\frac{\sigma}{\sqrt{db}}\;+\;\frac{16\eta}{\mu_{\mathcal{D}}}\,(L_{0}+L_{1}\sigma).$
@@
We choose the stepsize $\eta$ minimizing the right-hand side of Theorem C.5 subject to $\eta\leq\tfrac{\mu_{\mathcal{D}}}{16L_{1}}$, namely
$\eta\;=\;\min\!\left(\tfrac{\mu_{\mathcal{D}}}{16L_{1}},\,\sqrt{\tfrac{F_{0}}{(L_{0}+L_{1}\sigma)\,T}}\right).$ (26)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla f({\bm{x}}_{t})\|\;=\;\frac{1}{\mu_{\mathcal{D}}}\,\mathcal{O}\!\left(\frac{L_{1}F_{0}}{\mu_{\mathcal{D}}T}+\sqrt{\tfrac{(L_{0}+L_{1}\sigma)\,F_{0}}{T}}+\frac{\sigma}{\sqrt{d\,b}}\right).$ (27)
@@
For the uniform distribution over the unit sphere, we have $\mu_{\mathcal{D}}\asymp d^{-1/2}$. With $T=\Theta\!\big(dL_{1}/\varepsilon+d(L_{0}+L_{1}\sigma)F_{0}/\varepsilon^{2}\big)$ and $b=\Theta(\sigma^{2}/\varepsilon^{2})$, we ensure $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla f({\bm{x}}_{t})\|\leq\varepsilon$. The total number of function evaluations is then
$\textsc{Calls}=2bT\;=\;\mathcal{O}\!\left(\frac{d\,(L_{0}+L_{1}\sigma)\,F_{0}\,\sigma^{2}}{\varepsilon^{4}}\right).$
@@ Appendix D Finite sum case and Variance reduction
Under Assumption equation 28, suppose that $f$ is $(L_{0},L_{1})$-smooth, ${\bm{s}}_{t}\sim{\mathcal{D}}$ satisfies Assumption 2, and a variance-reduction scheme with snapshot period $m$ is used. Let $\eta_{t}\equiv\eta\leq\mu_{\mathcal{D}}/L_{1}$ and $F_{0}:=f({\bm{x}}_{0})-f^{\star}<\infty$. Then for each iteration, $\mathbb{E}[f({\bm{x}}_{t+1})]\;\leq\;\mathbb{E}[f({\bm{x}}_{t})]\;-\;\frac{\mu_{{\mathcal{D}}}}{2}\,\eta\,\mathbb{E}\|\nabla f({\bm{x}}_{t})\|\;+\;\frac{L_{0}}{2}\,\eta^{2}\;+\;\frac{4\,G\,m}{\sqrt{b}}.$ (29) Averaging over $t=0,\dots,T-1$, we obtain $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla f({\bm{x}}_{t})\|\;\leq\;\frac{1}{\mu_{{\mathcal{D}}}}\left(\frac{F_{0}}{\eta T}+\frac{L_{0}}{2}\,\eta+\frac{4\,G\,m}{\sqrt{b}}\right).$ (30)
@@ Appendix E Learning with Helper or Human Feedback: Convergence Proof
Suppose $f$ is $(L_{0},L_{1})$-smooth in the sense of equation 13, the random directions satisfy Assumption 2 ($\mathbb{E}\|{\bm{s}}_{t}\|^{2}=1$ and $\mathbb{E}[|\langle\nabla f({\bm{x}}),{\bm{s}}_{t}\rangle|]\geq\mu_{{\mathcal{D}}}\|\nabla f({\bm{x}})\|$), and the helper satisfies Assumption 7. Let $F_{0}:=f({\bm{x}}_{0})-f^{\star}<\infty$ and choose a constant stepsize $\eta_{t}\equiv\eta\leq\mu_{{\mathcal{D}}}/L_{1}$. Then $\mathbb{E}[f({\bm{x}}_{t+1})]\;\leq\;\mathbb{E}[f({\bm{x}}_{t})]\;-\;\frac{\mu_{{\mathcal{D}}}}{2}\,\eta\,\mathbb{E}\|\nabla f({\bm{x}}_{t})\|\;+\;\frac{L_{0}}{2}\,\eta^{2}\;+\;2\,\delta.$ (35) Averaging over $t=0,\dots,T-1$ yields $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla f({\bm{x}}_{t})\|\;\leq\;\frac{1}{\mu_{{\mathcal{D}}}}\left(\frac{F_{0}}{\eta T}+\frac{L_{0}}{2}\,\eta+\frac{2\delta}{\eta}\right).$ (36)
@@ Appendix F Failure of Classical Momenta
We now examine whether classical momentum techniques from first-order optimization can improve the performance of Random Search. We consider three natural extensions of popular momentum formulations: Heavy-Ball [polyak1964some], momentum-based variance reduction (MVR) [cutkosky2019momentum], and implicit gradient transport [arnold2019reducing]. Each method can be adapted to our comparison-based framework by replacing stochastic gradients with function differences.